OpenStudy (anonymous):

Confused... The question is: Identify the 13th term of the geometric sequence 10, 8, 6.4 My work: 10(.8)^12 = .6871947674 Doesn't match possible answers? o.O

OpenStudy (anonymous):

Possible answers:

OpenStudy (anonymous):

@nphuongsun93 This has me stumped..

OpenStudy (anonymous):

My class has not once gone over this.

OpenStudy (anonymous):

give me a bit, been a while since i do this

OpenStudy (anonymous):

^lol

OpenStudy (anonymous):

guess it's D, it's nearest to 0.69

OpenStudy (anonymous):

Alright, thanks. Got to say, this e-school just sucks.

OpenStudy (anonymous):

@nphuongsun93 I just found this answer + work online. What do you think? a(1) = 52 r = 39/52 = 29.25/39 = 3/4 = 0.75 a(n) = a(1) * r^(n - 1) a(15) = 52 * 0.75^(15 - 1) a(15) = 52 * 0.75^14 a(15) =~ 0.92653329670429229736328125

OpenStudy (anonymous):

yea it looks right o-o